# Explainable Prediction of the Mechanical Properties of Composites with CNNs

### v1

@@ Abstract
Composites are amongst the most important materials manufactured today, as evidenced by their use in countless applications. In order to establish the suitability of composites in specific applications, finite element (FE) modelling, a numerical method based on partial differential equations, is the industry standard for assessing their mechanical properties. However, FE modelling is exceptionally costly from a computational viewpoint, a limitation which has led to efforts towards applying AI models to this task. However, in these approaches: the chosen model architectures were rudimentary, feed-forward neural networks giving limited accuracy; the studies focus on predicting elastic mechanical properties, without considering material strength limits; and the models lacked transparency, hindering trustworthiness by users. In this paper, we show that convolutional neural networks (CNNs) equipped with methods from explainable AI (XAI) can be successfully deployed to solve this problem. Our approach uses customised CNNs trained on a dataset we generate using transverse tension tests in FE modelling to predict composites’ mechanical properties, i.e., Young’s modulus and yield strength. We show empirically that our approach achieves high accuracy, outperforming a baseline, ResNet-34, in estimating the mechanical properties. We then use SHAP and Integrated Gradients, two post-hoc XAI methods, to explain the predictions, showing that the CNNs use the critical geometrical features that influence the composites’ behaviour, thus allowing engineers to verify that the models are trustworthy by representing the science of composites.
@@ 1 Introduction
Composites are one of the most important materials that are manufactured today due to their exceptional specific mechanical properties, i.e., mechanical properties normalised by density. This has resulted in their widespread application, e.g. in wind turbines, aerospace components and sports equipment [19]. The evaluation of these mechanical properties, and thus the performance and suitability of composites in individual applications, can be assessed via a variety of analytical, experimental and numerical methods [9, 26, 27]. Analytical models provide accurate estimations for simplistic cases but rely on idealised assumptions (e.g. uniform inclusion positions and geometries) and do not account for all necessary geometrical or material parameters (e.g. weak bonding and complex constitutive laws) [7]. Experimental methods can be challenging and time-consuming [8], especially when analysing how and why a material or system fails. Further, post-failure analysis often requires multiple steps (e.g. cutting samples and using optical tools), while in-situ testing demands specialised and costly equipment (e.g. in-situ testing rigs, X-ray tomography equipment, or high-speed cameras) [29]. These requirements can make experimental methods less practical for studies that explore how different factors influence failure behaviour. Numerical models, e.g. employing finite element (FE) modelling, can provide accurate estimations of composites’ properties and incorporate more realistic assumptions about their geometries, unlike analytical models. However, they are limited to small, simulated volumes containing only a few inclusions due to their extensive computational cost [16, 6] (in some cases simulations required up to 13 days to complete [26]). Attempts have been made within the composite research community to develop more efficient numerical models, but these are often tailored to specific loading scenarios and types of composites and require prior knowledge of the key underlying physical mechanisms governing the composite behaviour [15, 3]. Thus, while numerical models are the favoured option for assessing the properties of composites, they are limited by these weaknesses.
@@
Concurrently, the inherent versatility of AI models in handling various types of data (e.g. tabular or images), along with their computational efficiency (especially after a one-off training cost), has been demonstrated by their widespread deployment in engineering problems [18, 22, 30]. In composite engineering, AI models have recently been applied to various tasks, including property prediction [7], sensitivity analysis of failure processes [17, 28], optimisation of composite properties [1] and manufacturing design [18]. Some efforts have been made to apply AI models to predicting mechanical properties in composites [7, 16, 33]. However, these studies either focus solely on predicting elastic properties [7] or, when extending to strength limits, are constrained by limited training data (approximately 1,000 datapoints) and small composite microstructures (fewer than 15 fibres) [16, 33]. These limitations do not meet the criteria for statistically representative composite microstructures [21], raising concerns about the reliability of the results. Furthermore, most of the deployed AI model architectures, primarily feed-forward neural networks, remain relatively rudimentary, resulting in limited accuracy [7, 16]. Finally, the AI models were not supplemented with explanations to allow engineers to verify that the patterns being learnt are indeed representing the science of composites, thus missing a crucial opportunity for fostering trust in the engineers.
@@
In this paper, we propose a solution to this problem by adapting convolutional neural networks (CNNs, [13]), supplemented with techniques from explainable AI (XAI, see [2] for a recent overview), to this image regression task. To ensure that we have a representative training dataset for our CNN model and that we overcome the reported computational cost associated with 3D FE modelling of composites, we propose an FE-based framework for 2D composite microstructures that contain at least 100 fibres (a criterion for representative composite microstructures [21]), as illustrated in Figure 1, to predict composite properties under transverse tension virtual tests. We generate a database of 2D composite microstructures, where the constituents exist in different proportions and spatial arrangements of fibres, along with estimates of their mechanical properties (Young’s modulus and yield strength), obtained through FE-based tensile tests. We then deploy XAI methods based on feature attribution, to explain the predictions, thus showing that the CNNs use the critical geometrical features that influence the composites’ behaviour. In doing so, we demonstrate that not only does our methodology produce predictions which align closely with FE modelling in this budding research area, it also allows engineers to verify that the CNNs are representing the science of composites, thus increasing its trustworthiness.
@@ 2 Related Work
Property Prediction.   Several studies have employed 2D FE modelling at the microscale to explicitly simulate the geometry of composite constituents (i.e., fibres, which provide the structural integrity of the composite, and a matrix, which holds the fibres in place) and investigate how variations in the spatial arrangement and proportion of fibres affect the composite’s performance under transverse loading. These studies have analysed fibres with irregular shapes [7], stochastic geometries (e.g. checkerboard arrangements [1]), and spherical shapes [7, 16, 33]. The output from these simulations served as training data for AI models, including artificial neural networks (ANNs, i.e., feed-forward multi-layer perceptrons), and CNNs, demonstrating high accuracy compared to FE simulations (treated as the ground truth here). However, these studies exhibit certain limitations. Some focus solely on predicting elastic moduli [7], properties that can already be accurately estimated using analytical models. Others require additional pre-processing steps, such as principal component analysis [7, 16], before 2D composite microstructures can be used in AI models, while others utilise a small, non-representative training dataset [16, 33].
@@
XAI in Engineering.   XAI can enhance interpretability by providing insights into AI model decisions and uncovering information that finite element models alone may not readily offer. Integrating AI with FE modelling in a synergistic way can enhance our understanding of the geometrical mechanisms influencing composite performance. However, despite its potential, XAI has not yet been widely applied in the composite engineering field. A key example demonstrating the importance of XAI in composite applications is the work by Yossef et al. [32] which introduced a powerful AI framework integrating FE modelling and machine learning models. In their study, FE modelling was used to predict composite strength (i.e., the FE modelling output) based on the material properties of laminates with different ply layup stacking sequences (i.e., the FE modelling input). A dataset of 1,000 composite strength values from the FE modelling was then employed to train an AI model (XGBoost [5]). XGBoost was then coupled with XAI techniques (including SHAP [14]) to identify the material properties with the greatest influence on composite strength, suggest an optimal laminate layup, and demonstrate the potential of XAI for advancing scientific discovery in this domain.
@@ 3.1 Microstructure Image Generation
A total of 7,893 2D composite microstructures  [^4] Only 7,893 out of 13,000 generated 2D microstructures were analysed. The remaining cases were excluded due to FE meshing failures, primarily from closely spaced fibres, and occasional disruptions in job scheduling on the high-performance computing system, which led to simulations with errors. were generated, with the fibre volume fraction, $V^{f}$, ranging from 20.0% to 44.0% in discrete intervals of 2.0%. Although the fibre volume fraction in composites can reach up to $V^{f}$= 60.0%, microstructures with $V^{f}$ > 44.0% were not considered due to limitations in the generation algorithm, a known constraint of RSA techniques. The 2D composite image on the left of the top row of Figure 1 shows a typical microstructure with a fibre volume fraction of 40.0%.
@@ 3.2 Calculation of the Mechanical Properties via FE Modelling
The final task of the FE modelling was to compute the bulk mechanical properties of the composite, specifically the elastic modulus ($E^{c}$) and yield strength ($\sigma_{y}^{c}$). These values were later used as the predicted variables in the training data for the CNNs. The FE modelling provided outputs in the form of reaction forces ($F_{zz}^{c}$) and displacements in the z-direction ($u_{zz}^{c}$) at the right vertical edge of the microstructure (z = $h^{c}$), as shown in the image on the second from the right of the top row of Figure 1. These outputs were then converted into stress and strain values representative of the composite’s macroscopic behaviour. The stress, $\sigma_{zz}^{c}$, was calculated by dividing the reaction force at the right edge ($F_{zz}^{c}$) by the original cross-sectional area of the composite, defined as $A^{c}=h^{c}$ (unit thickness). The corresponding strain, $\epsilon_{zz}^{c}$, was determined from the displacement of the right edge ($u_{zz}^{c}$) divided by the initial length of the composite microstructure, $L^{c}=h^{c}$. The stress–strain relationship could then be plotted as in the image on the right of the top row of Figure 1.
The Young’s modulus of the composite ($E^{c}$) was determined as the gradient of the $\sigma_{zz}^{c}$ $\epsilon_{zz}^{c}$ curve at the first increment of the simulation. The composite yield strength ($\sigma_{y}^{c}$) was approximated using the 0.2% proof stress convention. The image on the right of the top row of Figure 1 illustrates a typical example, highlighting the locations of $E^{c}$ and $\sigma_{y}^{c}$ measurements for a 2D composite microstructure with a $V^{f}=40.0\%$. The outputs for $E^{c}$ and $\sigma_{y}^{c}$ from the 7,893 FE modelling simulations were subsequently used for the CNN prediction (§4) and application of XAI techniques (§5).
@@ 3.3 Evaluation of the Mechanical Properties
Figure 3 summarises the results of the FE modelling of the 7,893 composite microstructures in terms of the Young’s modulus and the yield strength of the composite as a function of the fibre volume fraction. Both plots indicate a linear relationship between these two material properties and the fibre volume fraction, as evidenced by the high Pearson correlation factor, $r$. This linear relationship between the composite’s material properties and fibre volume fraction is well established and widely reported in the composite engineering field for fibre volume fraction values less than 50% [20]. While $V^{f}$ is widely reported as the dominant influencing factor in $E^{c}$ predictions [7, 16, 20], the slightly greater variability observed in the $\sigma_{y}^{c}$ vs. $V^{f}$ trend suggests that $\sigma_{y}^{c}$ may be influenced by additional geometrical parameters, such as local fibre arrangements [16] and interfibre distances. The XAI analysis presented in §5 will shed further insights into the contribution of the spatial fibre arrangements to materials properties of the composite.
@@ 4.1 Model Architectures and Training
Our baseline architecture was an instance of ResNet-34, given its excellent performance across a range of tasks [10]. The off-the-shelf ResNet-34 architecture was modified such that the final layer adjusted to contain one node to allow for a regression output. We then implemented a customised CNN of the architecture shown in Figure 4, which was the result of extensive hyperparameter optimisation.
@@ 4.2 Prediction Results
Sample (10%) ResNet-34 $-1.2205$ $58$ $-3.3193$ $61$
Sample (10%) Customised CNN $0.9583$ $5$ $0.9319$ $5$
@@
Table 2 shows our results for the CNNs in the form of their correlation with FE modelling and training time in predicting the mechanical properties on the sample and full datasets. It can be seen that the ResNet-34 architecture did not perform well in predicting the mechanical properties, with $R^{2}$ values of $-1.2205$ for $E^{c}$ and $-3.3193$ for $\sigma_{y}^{c}$. We believe that ResNet-34 would have likely achieved significantly better correlation with more training, but that this architecture is somewhat overkill for this relatively simple task, given that our (much simpler) customised CNN performs well even when trained on the sample dataset. This is backed up by the training times (also shown in Table 2), with our CNN being an order of magnitude faster.
@@ 5 Explanation of the CNN Predictions
We now assess, with a focus on the science of composites, the explanations from two post-hoc XAI methods, SHAP [14] and Integrated Gradients (IG) [24], for the predictions made by our customised CNN models.
@@ 6 Conclusions and Future Work
Our findings open up many new opportunities in the composite engineering field. Specifically, these advancements enable improvements in property prediction through various mechanisms including the following:
@@
  • Another key development would involve accounting for the multiphase nature of composites by explicitly simulating the interface between the fibres and the matrix. The bonding at this interface is often imperfect, which can significantly influence composite performance and lead to failure mechanisms such as fibre-matrix debonding. Capturing this interfacial effect, especially when combined with the use of advanced constitutive laws (as mentioned above), would provide a more comprehensive and accurate framework for predicting the mechanical behaviour of composites.
  • We plan to expand our approach to 3D FE modelling and other loading scenarios, such as longitudinal compression, tension, and shear. Composites exhibit highly anisotropic behaviour, meaning their mechanical response varies significantly depending on the loading direction and fibre orientation. As stated in §1, 3D FE modelling of composites is computationally expensive for parametric investigations. Thus, an accurate AI model would prove a valuable tool in multiscale composite modelling (micro–meso–macro), a challenge that has drawn attention in composite research for decades.
@@
Bikos kindly acknowledges additional funding for this research provided by the UK Engineering and Physical Sciences Research Council (EPSRC) programme Grant EP/T011653/1, Next Generation Fibre-Reinforced Composites (NextCOMP): a Full-Scale Redesign for Compression in a collaboration between Imperial College London and the University of Bristol. Rago and Toni were partially funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101020934) and by J.P. Morgan and the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme (grant agreement no. RCSRF2021\11\45). For the purpose of open access, a Creative Commons Attribution (CC BY-NC 4.0) license is applied to any author-accepted manuscript version arising.

### v2

@@ Abstract
Composites are amongst the most important materials manufactured today, as evidenced by their use in countless applications. In order to establish the suitability of composites in specific applications, finite element (FE) modelling, a numerical method based on partial differential equations, is the industry standard for assessing their mechanical properties. However, FE modelling is exceptionally costly from a computational viewpoint, a limitation which has led to efforts towards applying AI models to this task. However, in these approaches: the chosen model architectures were rudimentary, feed-forward neural networks giving limited accuracy; the studies focused on predicting elastic mechanical properties, without considering material strength limits; and the models lacked transparency, hindering trustworthiness by users. In this paper, we show that convolutional neural networks (CNNs) equipped with methods from explainable AI (XAI) can be successfully deployed to solve this problem. Our approach uses customised CNNs trained on a dataset we generate using transverse tension tests in FE modelling to predict composites’ mechanical properties, i.e., Young’s modulus and yield strength. We show empirically that our approach achieves high accuracy, outperforming a baseline, ResNet-34, in estimating the mechanical properties. We then use SHAP and Integrated Gradients, two post-hoc XAI methods, to explain the predictions, showing that the CNNs use the critical geometrical features that influence the composites’ behaviour, thus allowing engineers to verify that the models are trustworthy by representing the science of composites.
@@ 1 Introduction
Composites are one of the most important materials that are manufactured today due to their exceptional specific mechanical properties, i.e., mechanical properties normalised by density. This has resulted in their widespread application, e.g. in wind turbines, aerospace components and sports equipment [18]. The evaluation of these mechanical properties, and thus the performance and suitability of composites in individual applications, can be assessed via a variety of analytical, experimental and numerical methods [8, 25]. Analytical models provide accurate estimations for simplistic cases but rely on idealised assumptions (e.g. uniform inclusion positions and geometries) and do not account for all necessary geometrical or material parameters (e.g. weak bonding and complex constitutive laws) [6]. Experimental methods can be challenging and time-consuming [7], especially when analysing how and why a material or system fails. Furthermore, post-failure analysis often requires multiple steps (e.g. cutting samples and using optical tools), while in-situ testing demands specialised and costly equipment (e.g. in-situ testing rigs, X-ray tomography equipment, or high-speed cameras) [27]. These requirements can make experimental methods less practical for studies that explore how different factors influence failure behaviour. Numerical models, e.g. employing finite element (FE) modelling, can provide accurate estimations of composites’ properties and incorporate more realistic assumptions about their geometries, unlike analytical models. However, they are limited to small, simulated volumes containing only a few inclusions due to their extensive computational cost [15] (in some cases simulations required up to 13 days to complete [25]). Attempts have been made within the composite research community to develop more efficient numerical models, but these are often tailored to specific loading scenarios and types of composites and require prior knowledge of the key underlying physical mechanisms governing the composite behaviour [14, 3]. Thus, while numerical models are the favoured option for assessing the properties of composites, they are limited by these weaknesses.
@@
Concurrently, the inherent versatility of AI models in handling various types of data (e.g. tabular or images), along with their computational efficiency (especially after a one-off training cost), has been demonstrated by their widespread deployment in engineering problems [17, 21, 28]. In composite engineering, AI models have recently been applied to various tasks, including property prediction [6], sensitivity analysis of failure processes [16, 26], optimisation of composite properties [1] and manufacturing design [17]. Some efforts have been made to apply AI models to predicting mechanical properties in composites [6, 15, 31]. However, these studies either focus solely on predicting elastic properties [6] or, when extending to strength limits, are constrained by limited training data (approximately 1,000 datapoints) and small composite microstructures (fewer than 15 fibres) [15, 31]. These limitations do not meet the criteria for statistically representative composite microstructures [20], raising concerns about the reliability of the results. Furthermore, most of the deployed AI model architectures, primarily feed-forward neural networks, remain relatively rudimentary, resulting in limited accuracy [6, 15]. Finally, these AI models were not supplemented with explanations to allow engineers to verify that the patterns being learnt are indeed representing the science of composites, thus missing a crucial opportunity for fostering trust in the engineers.
@@
In this paper, we propose a solution to this problem by adapting convolutional neural networks (CNNs) [12], supplemented with techniques from explainable AI (XAI, see [2] for a recent overview), to this image regression task. To ensure that we have a representative training dataset for our CNN model and that we overcome the reported computational cost associated with 3D FE modelling of composites, we propose an FE-based framework for 2D composite microstructures that contain at least 100 fibres (a criterion for representative composite microstructures [20]), as illustrated in Figure 1, to predict composite properties under transverse tension virtual tests. We generate a database of 2D composite microstructures, where the constituents exist in different proportions and spatial arrangements of fibres, along with estimates of their mechanical properties (Young’s modulus and yield strength), obtained through FE-based tensile tests. We then deploy XAI methods based on feature attribution, to explain the predictions, thus showing that the CNNs use the critical geometrical features that influence the composites’ behaviour. In doing so, we demonstrate that not only does our methodology produce predictions which align closely with FE modelling in this budding research area, it also allows engineers to verify that the CNNs are representing the science of composites, thus increasing its trustworthiness.
@@ 2 Related Work
Property Prediction.   Several studies have employed 2D FE modelling at the microscale to explicitly simulate the geometry of composite constituents (i.e., fibres, which provide the structural integrity of the composite, and a matrix, which holds the fibres in place) and investigate how variations in the spatial arrangement and proportion of fibres affect the composite’s performance under transverse loading. These studies have analysed fibres with irregular shapes [6], stochastic geometries (e.g. checkerboard arrangements [1]), and spherical shapes [6, 15, 31]. The outputs from these simulations served as training data for AI models, including artificial neural networks (ANNs, i.e., feed-forward multi-layer perceptrons), and CNNs, demonstrating high accuracy compared to FE simulations (treated as the ground truth here). However, these studies exhibit certain limitations. Some focus solely on predicting elastic moduli [6], properties that can already be accurately estimated using analytical models. Others require additional pre-processing steps, such as principal component analysis [6, 15], before 2D composite microstructures can be used in AI models, while others utilise a small, non-representative training dataset [15, 31].
@@
XAI in Engineering.   XAI can enhance interpretability by providing insights into AI model decisions and uncovering information that finite element models alone may not readily offer. Integrating AI with FE modelling in a synergistic way can enhance our understanding of the geometrical mechanisms influencing composite performance. However, despite its potential, XAI has not yet been widely applied in the composite engineering field. A key example demonstrating the importance of XAI in composite applications is the work by Yossef et al. [30] which introduced a powerful AI framework integrating FE modelling and machine learning models. In their study, FE modelling was used to predict composite strength (i.e., the FE modelling output) based on the material properties of laminates with different ply stacking sequences (i.e., the FE modelling input). A dataset of 1,000 composite strength values from the FE modelling was then employed to train an AI model (XGBoost [5]). XGBoost was then coupled with XAI techniques (including SHAP [13]) to identify the material properties with the greatest influence on composite strength, suggest an optimal laminate layup, and demonstrate the potential of XAI for advancing scientific discovery in this domain.
@@ 3.1 Microstructure Image Generation
A total of 7,893 2D composite microstructures were generated [^4] Only 7,893 out of 13,000 generated 2D microstructures were analysed. The remaining cases were excluded due to FE meshing failures, primarily from closely spaced fibres, and occasional disruptions in job scheduling on the high-performance computing system, which led to simulations with errors. , with the fibre volume fraction, $V^{f}$, ranging from 20.0% to 44.0% in discrete intervals of 2.0%. Although the fibre volume fraction in composites can reach up to $V^{f}$= 60.0%, microstructures with $V^{f}$ > 44.0% were not considered due to limitations in the generation algorithm, a known constraint of RSA techniques. The 2D composite image on the left of the top row of Figure 1 shows a typical microstructure with a fibre volume fraction of 40.0%.
@@ 3.2 Calculation of the Mechanical Properties via FE Modelling
The final task of the FE modelling was to compute the bulk mechanical properties of the composite, specifically the elastic modulus ($E^{c}$) and yield strength ($\sigma_{y}^{c}$). These values were later used as the predicted variables in the training data for the CNNs. The FE modelling provided outputs in the form of reaction forces ($F_{zz}^{c}$) and displacements in the z-direction ($u_{zz}^{c}$) at the right vertical edge of the microstructure (z = $h^{c}$), as shown in the image on the second from the right of the top row of Figure 1. These outputs were then converted into stress and strain values representative of the composite’s macroscopic behaviour. The stress, $\sigma_{zz}^{c}$, was calculated by dividing the reaction force at the right edge ($F_{zz}^{c}$) by the original cross-sectional area of the composite, defined as $A^{c}=h^{c}$ × (unit thickness). The corresponding strain, $\epsilon_{zz}^{c}$, was determined from the displacement of the right edge ($u_{zz}^{c}$) divided by the initial length of the composite microstructure, $L^{c}=h^{c}$. The stress–strain relationship could then be plotted as in the image on the right of the top row of Figure 1.
The Young’s modulus of the composite ($E^{c}$) was determined as the gradient of the $\sigma_{zz}^{c}$ - $\epsilon_{zz}^{c}$ curve at the first increment of the simulation. The composite yield strength ($\sigma_{y}^{c}$) was approximated using the 0.2% proof stress convention. The image on the right of the top row of Figure 1 illustrates a typical example, highlighting the locations of $E^{c}$ and $\sigma_{y}^{c}$ measurements for a 2D composite microstructure with a $V^{f}=40.0\%$. The outputs for $E^{c}$ and $\sigma_{y}^{c}$ from the 7,893 FE modelling simulations were subsequently used for the CNN prediction (§4) and application of XAI techniques (§5).
@@ 3.3 Evaluation of the Mechanical Properties
Figure 3 summarises the results of the FE modelling of the 7,893 composite microstructures in terms of the Young’s modulus and the yield strength of the composite as a function of the fibre volume fraction. Both plots indicate a linear relationship between these two material properties and the fibre volume fraction, as evidenced by the high Pearson correlation factor, $r$. This linear relationship between the composite’s material properties and fibre volume fraction is well established and widely reported in the composite engineering field for fibre volume fraction values less than 50% [19]. While $V^{f}$ is widely reported as the dominant influencing factor in $E^{c}$ predictions [6, 15, 19], the slightly greater variability observed in the $\sigma_{y}^{c}$ vs. $V^{f}$ trend suggests that $\sigma_{y}^{c}$ may be influenced by additional geometrical parameters, such as local fibre arrangements [15] and interfibre distances. The XAI analysis presented in §5 will shed further light on the contribution of the spatial fibre arrangements to materials properties of the composite.
@@ 4.1 Model Architectures and Training
Our baseline architecture was ResNet-34, given its excellent performance history [9], modified such that the final layer adjusted to contain one node to allow for a regression output. We then implemented our customised CNN of the architecture shown in Figure 4, which was the result of extensive hyperparameter optimisation. (We leave the testing of other configurations of CNNs to future work.)
@@ 4.2 Prediction Results
Sample (10%) ResNet-34 1.2205-1.2205 $58$ 3.3193-3.3193 $61$
Sample (10%) Customised CNN $0.9583$ 55 $0.9319$ 55
@@
Table 2 shows our results for the CNNs in the form of their correlation with FE modelling and training time in predicting the mechanical properties on the sample and full datasets. It can be seen that the ResNet-34 architecture did not perform well in predicting the mechanical properties, with $R^{2}$ values of 1.2205-1.2205 for $E^{c}$ and 3.3193-3.3193 for $\sigma_{y}^{c}$. We believe that ResNet-34 would have likely achieved significantly better correlation with more training, but this architecture is somewhat overkill for this relatively simple task, given that our (much simpler) customised CNN performs well even when trained on the sample dataset. This is backed up by the training times (also shown in Table 2), with our CNN being an order of magnitude faster.
@@ 5 Explanation of the CNN Predictions
We now assess, with a focus on the science of composites, the explanations from two post-hoc XAI methods, SHAP [13] and Integrated Gradients (IG) [23], for the predictions made by our customised CNN models. SHAP was implemented using in-built functions from the SHAP Python library [13], whilst IG was implemented using the image visualisation method [23].
@@ 6 Conclusions and Future Work
Our findings open up many new opportunities in the composite engineering field, including the following:
  • Assessing the predictive power of our trained CNNs on real composite microstructures from micro-Computed Tomography tests, which exhibit higher fibre content than the synthetic data, fibre–fibre contact, and local variations in fibre distribution (fibre- or matrix-rich regions).
@@
  • Accounting for the multiphase nature of composites by explicitly simulating the interface between the fibres and the matrix. The bonding at this interface is often imperfect, which can significantly influence composite performance and lead to failure mechanisms such as fibre-matrix debonding. Capturing this interfacial effect, especially when combined with the use of advanced constitutive laws (as mentioned above), would provide a more comprehensive and accurate framework for predicting the mechanical behaviour of composites.
  • Expanding our approach to 3D FE modelling and other loading scenarios, such as longitudinal compression, tension, and shear. Composites exhibit highly anisotropic behaviour, meaning their mechanical response varies significantly depending on the loading direction and fibre orientation. As stated in §1, 3D FE modelling of composites is computationally expensive for parametric investigations. Thus, an accurate AI model would prove a valuable tool in multiscale composite modelling (micro–meso–macro), a challenge that has drawn attention in composite research for decades.
@@
Bikos kindly acknowledges additional funding for this research provided by the UK Engineering and Physical Sciences Research Council (EPSRC) programme Grant EP/T011653/1, Next Generation Fibre-Reinforced Composites (NextCOMP): a Full-Scale Redesign for Compression in a collaboration between Imperial College London and the University of Bristol. Rago and Toni were partially funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101020934) and by J.P. Morgan and the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme (grant agreement no. RCSRF2021\11\45). A Creative Commons Attribution (CC BY-NC 4.0) license is applied to any author-accepted manuscript version arising.